Find the solution of Non-exact D.E


Given:


(x+4y^3)dy-ydx=0

1
Expert's answer
2021-12-12T17:54:17-0500

Let us find the solution of differential equation "(x+4y^3)dy-ydx=0." It follows that "y=0" is a solution. If "y\ne 0," then the last equation is equivalent to "x+4y^3-yx'=0," and hence to "yx'-x=4y^3." Let us divide both parts by "y^2." Then we get the differential equation "\frac{1}yx'-\frac{1}{y^2}x=4y," which is equivalent to "(\frac{1}yx)'=4y." It follows that "\frac{1}yx=2y^2+C." We conclude that the general solution of the differential equation is

"x=2y^3+Cy,\ y=0."